OpenStudy (anonymous):

If the endpoints of a diameter of a circle are A(2, 5) and B (7, 2), what is the radius of the circle?

OpenStudy (anonymous):

You would do Midpoint first and then distance formulas right?

OpenStudy (anonymous):

@MathMan321

OpenStudy (science0229):



OpenStudy (science0229):

\[2r=\sqrt{(7-2)^2+(2-5)^2}\] \[2r=\sqrt{34}\] So the radisu is sqrt(34)/2

OpenStudy (anonymous):

so the answer would be Sqrt(17)

OpenStudy (science0229):

Nope. \[\frac{ \sqrt{34} }{ 2 }=\frac{ \sqrt{34} }{ \sqrt{4} }=\sqrt{\frac{ 34 }{ 4 }}=\sqrt\frac{ 17 }{ 2 }\]

OpenStudy (anonymous):

So it would be sqrt(17/2)

OpenStudy (anonymous):

why did you go from 2 to sqrt of 4

OpenStudy (science0229):

There is a property of square root\[\frac{ \sqrt{a} }{ \sqrt{b} }=\sqrt{\frac{ a }{ b }}\] where a>=0 and b>0

OpenStudy (anonymous):

oh okay, so the answer wuld be sqrt17/2

OpenStudy (science0229):

yep

OpenStudy (anonymous):

thanks

OpenStudy (science0229):

welcome :)

OpenStudy (anonymous):

Can you help me with this one it is a true or false question I just meed feedback.

OpenStudy (anonymous):

@science0229

OpenStudy (science0229):

yeah